Sandia Corporation manufactures metal toolboxes. It adds all materials at the beginning of the manufacturing process. The company has provided the following information:

Units Costs
Beginning work in process (23% complete) 31,000
Direct materials $ 42,000
Conversion cost 90,000
Total cost of beginning work in process $ 132,000
Number of units started 73,000
Number of units completed and transferred to finished goods ?
Ending work in process (59% complete) 90,000
Current period costs
Direct materials $ 80,000
Conversion cost 157,000
Total current period costs $ 237,000

Required:

1. Using the weighted-average method of process costing, complete each of the following steps:

a. Reconcile the number of physical units worked on during the period.

Physical Units Physical Units
Beginning Units Units Completed
Units Started Ending Units
Total Units 0 Total Units 0

b. Calculate the number of equivalent units.

Equivalent Units
Physical Units Direct Materials Conversion
Units Completed
Ending Inventory
Total 0 0 0

c. Calculate the cost per equivalent unit.

Direct Materials Conversion
Cost per Equivalent Unit

d. Reconcile the total cost of work in process.

Direct Materials Conversion Total Cost
Units Completed
Ending Inventory
Total Cost Accounted for $0 $0 $0
Answer #1

a. Reconcile the number of physical units worked on during the period.

Physical Units Physical Units
Beginning Units 31000 Units Completed 14000
Units Started 73000 Ending Units 90000
Total Units 104000 Total Units 104000

b. Calculate the number of equivalent units.

Equivalent Units
Physical Units Direct Materials Conversion
Units Completed 14000 14000 14000
Ending Inventory 90000 90000 90000*59% = 53100
Total 104000 104000 67100

c. Calculate the cost per equivalent unit.

Direct Materials Conversion
Cost per Equivalent Unit (42000+80000)/104000 =1.173 247000/67100 = 3.681

d. Reconcile the total cost of work in process.

Direct Materials Conversion Total Cost
Units Completed 16422 51534 67956
Ending Inventory 105570 195461 301031
Total Cost Accounted for 121992 246995 368987
Answer #2

How did you get the last table? letter d

answered by: oriana